OpenStudy (anonymous):

how do you find the sides of a triangle if you only have the longest side found

OpenStudy (saifoo.khan):

if it's a right angle triangle, use Pythagoras.

OpenStudy (anonymous):

use sin cos tan

OpenStudy (anonymous):

i dont know the angle only the side

hero (hero):

Draw the triangle and label it for all of us to see

hero (hero):

 Yay! I created a triangle using that drawing feature below. Now you can do the same :D

OpenStudy (pokemon23):

omg openstudy work?

hero (hero):

Shocking, but true

OpenStudy (pokemon23):

a^2+b^2=c^2

OpenStudy (anonymous):

i know that equation

OpenStudy (pokemon23):

very nice :)

hero (hero):

What about drawing the triangle?

OpenStudy (pokemon23):

triangle help :)

OpenStudy (anonymous):

i have the picture of the triangle but its a bigger problem

hero (hero):

Can you try drawing it?

OpenStudy (anonymous):

and are us till looking for a side measurement ?

OpenStudy (anonymous):

yes ive tried the pythag and sin cos tan and its still the wrong answers

hero (hero):

I see why you didn't draw it

OpenStudy (anonymous):

i can try i sent it in an attachment

OpenStudy (anonymous):

umm your going to have to solve it with the circles...like looking for circumference and ratio

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

ur trying to find from A to B right?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

o well it gives u what yo need eb=8.4 andFp=3.2 so the way i see it just do 8.4*2 and 3.2*2 and add them

OpenStudy (anonymous):

ill try it

OpenStudy (anonymous):

and that gives u one more side then u should be able to use pathagereom

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

that was the wrong answer

OpenStudy (anonymous):

did u use a^2+B^2=c^2?

OpenStudy (anonymous):

yup

OpenStudy (anonymous):

hmm give me a sec

OpenStudy (anonymous):

 is this how u did it?